# Homotopy Methods for Convex Optimization

### v1

@@ Abstract
Convex optimization encompasses a wide range of optimization problems, containing many efficiently solvable subclasses. Interior point methods are currently the state-of-the-art approach for solving such problems, particularly effective for classes like semidefinite programming, quadratic programming, and geometric programming. However, their success hinges on the construction of self-concordant barrier functions for the feasible sets. In this work, we introduce an alternative method for tackling convex optimization problems, employing a homotopy. With this technique, the feasible set of a trivial optimization problem is continuously transformed into the target one, while tracking the solutions. We conduct an analysis of this approach, focusing on its application to semidefinite programs, hyperbolic programs, and convex optimization problems with a single convexity constraint. Moreover, we demonstrate that our approach numerically outperforms state-of-the-art methods in several interesting cases.
@@ 1 Introduction
Optimization problems with convex constraints and a convex objective function (aka convex optimization problems) are a class of optimization problems that find applications in quantum information theory, portfolio optimization, data fitting, and many more [5]. The state-of-the-art methods to solve these problems are interior-point methods, that go back to Nesterov and Nemirovski [21]. While these are successfully applied to many sub-classes of convex optimization problems, including semidefinite programming, quadratically constrained quadratic programs, or geometric programming, they don’t solve arbitrary convex optimization problems efficiently. The success of these methods heavily rely on the construction of a self-concordant barrier function.
In this paper, we present a way to solve certain convex optimization problems using a homotopy method. We apply this method to convex optimization problems with a single convexity constraint, as well as semidefinite and hyperbolic programs. The presented method, as all homotopy methods, is based on the following general idea:
@@
More precisely, we start from a convex optimization problem with an obvious solution (i.e. analytically solvable) and homotopically change it to the target problem (see Fig. 1). In contrast to some other homotopy methods, we propose to change the feasible set of the optimization problem, instead of changing the objective function. We then show that the path of optimal solutions is the unique solution of a system of differential equations, and we numerically solve this system to obtain a solution of the optimization problem.
@@
Main contributions.
We introduce the notion of a smooth and regular description of families of convex sets, and show that the path of optimal solutions is the unique solution of a certain differential equation (see Theorem 6 for the main result of this paper). We then construct explicit homotopies for two optimization classes and show that these homotopies are smooth and regular. These examples include:
@@
Homotopic approaches in other fields.
The idea of homotopy methods has been successfully applied in various different fields, including algebraic geometry, non-linear partial differential equations, and quantum computing:
  • $\triangleright$
    In numerical algebraic geometry (originally developed by Sommese and Wampler [31], see [11, 32] for an overview) the main objective is to solve a system of polynomial equations
    $p_{1}(\mathbf{z})=0,\ldots,p_{k}(\mathbf{z})=0$
    in multiple variables $\mathbf{z}\in\mathbb{C}^{n}$ with finitely many solutions. First, a system of polynomial equations
    $g_{1}(\mathbf{z})=0,\ldots,g_{k}(\mathbf{z})=0$
    with trivial solutions is generated. This could be for example the unit polynomials $g_{i}(\mathbf{z})=z_{i}^{d}$. Afterwards, these solutions are tracked via the homotopy
    $(1-t)g_{i}(\mathbf{z})+tp_{i}(\mathbf{z})=0\text{ for every }i\in\{1,\ldots,k\}.$
    More specifically, the homotopy method consists of letting evolve the parameter $t$ from 00 to $1$ and following the path of solutions along $t$ via solving a differential equation.
  • $\triangleright$
    The homotopy analysis method [17, 19, 18] and the homotopy perturbation method [13] are approaches to solve non-linear partial differential equation via infinitely many linear ones. By setting up a homotopy between a linear partial differential operator $\mathcal{L}$ to a non-linear one $\mathcal{N}$, one obtains a homotopy of solutions $(u_{t})_{t\in[0,1]}$, where $u_{0}$ is the solution of the initial linear equation $\mathcal{L}$ and $u_{1}$ is the solution of the target non-linear equation $\mathcal{N}$. Via a Taylor expansion
    $u_{t}(\mathbf{x})=u_{0}(\mathbf{x})+\sum_{k=1}^{\infty}v_{k}(\mathbf{x})t^{k}$
    one obtains linear differential equations for every $v_{k}$.
  • $\triangleright$
    In adiabatic quantum computing and quantum annealing [9], the main objective is to compute the ground state of a quantum-Hamiltonian $H_{B}$. This is in general an $\mathsf{NP}$-hard optimization problem [3]. Let
    $H_{t}\coloneqq(1-t)\cdot H_{A}+t\cdot H_{B}$
    describe a physical system which is described at time $t=0$ by a Hamiltonian $H_{A}$ with a trivial ground state and at time $t=1$ by the target Hamiltonian $H_{B}$. If the evolution from $H_{A}$ to $H_{B}$ happens slowly enough, then it is guaranteed that the state of the system remains in its ground state for every $t,$ which implies that one can obtain the ground state of $H_{B}$ when knowing the ground state of $H_{A}$. Since ground-states of quantum Hamiltonians can embed $\mathsf{NP}$-complete optimization problems, adiabatic quantum computing can be employed as a homotopy method solving $\mathsf{NP}$-complete optimization problems [20].
Related work.
So far, homotopy methods have also been applied in semidefinite programming in different contexts, which are however different from what we propose in this paper:
  • $\triangleright$
    To apply the interior point method to semidefinite programs, they have to be strictly feasible. By introducing a perturbation on the feasible set, this condition can be relaxed [12, 15].
  • $\triangleright$
    Convex optimization problems with algebraic constraints can be solved via a system of polynomial equations arising from the Karush–Kuhn–Tucker (KKT) conditions (see for example [5]). This system can be solved via numerical algebraic geometry (see for example [6] for an implementation and [28] for a related problem).
Overview.
This paper is structured as follows: In Section 2 we introduce the notion of a homotopy between optimization problems, and then show under which conditions the induced differential equation of optimal solutions has a unique solution. In Section 3 we apply our homotopy method to two optimization classes: hyperbolic programs (which include semidefinite programs) and convex optimization problems with a single constraint function. In Section 4 we provide some numerical examples that compare our method to existing methods. In Section 5 we present open questions and further directions. Finally, we present in Appendix A another canonical homotopy between semidefinite programs which does not behave well regarding our method, and in Appendix B we show that the method also generalizes to quasi-concave objective functions $f$ and when introducing a homotopy on $f$. Appendix C contains the proof of boundedness of rigidly convex sets defined by elementary symmetric polynomials, used in Section 4.
@@ 2.1 Homotopies of convex sets and problem statement
We do this by continuously deforming the feasible set from an easy one into $C$, and keeping track of an optimal solution of the corresponding optimization problems along the way. Note that this construction can be generalized to concave functions $f$, by introducing a homotopy of objective functions. In Appendix B we elaborate on this more general case.
To be able to define and work with the homotopies, we will assume that the feasible sets come with a description as follows:
@@
such that $p_{t}$ provides a smooth description of $C_{t}$ for all $t\in[0,1]$.
@@
But under a certain continuity condition on the family $\left(C_{t}\right)_{t\in[0,1]}$, every homotopy of smooth descriptions in indeed regular. Assume, for example, that
$B:=\bigcup_{t\in[0,1]}\partial C_{t}\times\{t\}\subseteq\mathbb{R}^{n+1}$
is an $n$-dimensional manifold (with boundary $\partial C_{0}\cup\partial C_{1}$). Then at each point $(\mathbf{x},t)\in B$ the gradient of $p$ is nonzero, already in space direction, since $p_{t}$ is a smooth description of $C_{t}$. Thus the implicit function theorem implies that the zero set of $p$ is the graph of a function in $t$ and $n-1$ of the space variables, locally around $(\mathbf{x},t)$. But since $B$ is an $n$-dimensional manifold contained in this zero set/graph, it coincides with it locally.
@@ 2.2 The Differential Equation for Stationary Points
    $w_{i}=0$. Then $v_{i}w_{k}=0$. Since $w_{k}\neq 0$, we have that $v_{i}=0$ which implies that $v_{i}=\lambda w_{i}$.∎
@@ 2.3 Existence of a Unique Path of Optimal Solutions
Let $\left(C_{t}\right)_{t\in[0,1]}$ be a family of compact convex sets, all with nonempty interior in $\mathbb{R}^{n}$. Let $p$ be a homotopy of smooth descriptions satisfying the following properties:
@@ 3.1 Semidefinite and Hyperbolic Programming
Generalizing spectrahedra to rigidly convex sets (see Definition 7) leads to the notion of hyperbolic programming. Rigidly convex sets arise from so-called real zero polynomials and share many properties with spectrahedra. The generalized Lax conjecture actually postulates that every rigidly convex set is a spectrahedron. While this statement is true for $n=2$, it is open in higher dimensions [16].
Note that our homotopy approach is useful beyond the established methods for solving semidefinite programs. First, as just mentioned, it is not clear whether each rigidly convex set is actually a spectrahedron, so we solve a potentially larger class of problems. It seems that no efficient general methods to solve hyperbolic programs are available yet. Second, even if a set is known to be a spectrahedron, it might be hard to find an explicit linear matrix inequality defining it. Third, even if the linear matrix inequality is explicitly known, it might be of much larger size than the degree of the corresponding real zero polynomial [29, 27], whereas we work directly with this polynomial.
@@ 3.1.1 Rigidly Convex Sets
In the following, we define and examine the notion of rigidly convex sets. Moreover, we review a procedure to approximate every rigidly convex set by a smooth version thereof. Most of the content in this subsection is also part of [22, Section 2].
@@
Let $A(\mathbf{x})=I_{s}+x_{1}A_{1}+\cdots+x_{n}A_{n}$ be a linear matrix polynomial, and define the polynomial
@@ 3.1.2 Smoothing of Real Zero Polynomials
We now introduce the smoothing operator $\mathcal{F}_{\varepsilon}$ on the space of real-zero polynomials, that converts every rigidly convex set into a smooth rigidly convex set that approximates the original set. For this purpose, we introduce the Renegar derivative.
@@
Definition 11 (Smoothing of real zero polynomials).
Let $p\colon\mathbb{R}^{n}\to\mathbb{R}$ be a real zero polynomial. We define
$\mathcal{F}_{\varepsilon}[p]\coloneqq p+\varepsilon\partial_{R}p$
as the $\varepsilon$-smoothed version of $p$.
@@
    If $\lambda$ is a root of multiplicity $k\geqslant 2$ of $p_{\mathbf{x}}$, then $\lambda$ is a root of multiplicity $k-1$ of $\mathcal{F}_{\varepsilon}[p]_{\mathbf{x}}$. All other roots of $\mathcal{F}_{\varepsilon}[p]_{\mathbf{x}}$ have with multiplicity $1$.

### v2

@@ Abstract
Convex optimization encompasses a wide range of optimization problems that contain many efficiently solvable subclasses. Interior point methods are currently the state-of-the-art approach for solving such problems, particularly effective for classes like semidefinite programming, quadratic programming, and geometric programming. However, their success hinges on the construction of self-concordant barrier functions for feasible sets.
In this work, we investigate and develop a homotopy-based approach to solve convex optimization problems. While homotopy methods have been considered in optimization before, their potential for general convex programs remains underexplored. This approach gradually transforms the feasible set of a trivial optimization problem into the target one while tracking solutions by solving a differential equation, in contrast to traditional central path methods. We establish a criterion that ensures that the homotopy method correctly solves the optimization problem and prove the existence of such homotopies for several important classes, including semidefinite and hyperbolic programs. Furthermore, we demonstrate that our approach numerically outperforms state-of-the-art methods in hyperbolic programming, highlighting its practical advantages.
@@ 1 Introduction
Optimization problems with convex constraints and a convex objective function (aka convex optimization problems) are a class of optimization problems that find applications in quantum information theory, portfolio optimization, data fitting, and many more [4]. The state-of-the-art methods to solve these problems are interior-point methods, which go back to Nesterov and Nemirovski [18] (see also [17, 16, 31, 26] for detailed overviews). While these are successfully applied to many sub-classes of convex optimization problems, including semidefinite programming, quadratically constrained quadratic programs, or geometric programming, they don’t solve arbitrary convex optimization problems efficiently. The success of these methods heavily rely on the construction of a self-concordant barrier function.
In this paper, we investigate and develop a way to solve certain convex optimization problems using a homotopy method. We apply this method to convex optimization problems with a single convexity constraint, as well as semidefinite and hyperbolic programs. The presented method, as all homotopy methods, is based on the following general idea:
@@
More precisely, we start from a convex optimization problem with an obvious solution (i.e. analytically solvable) and homotopically change it to the target problem (see Fig. 1). In contrast to some other homotopy methods, we propose to change the feasible set of the optimization problem, instead of changing the objective function. We then show that the path of optimal solutions is the unique solution of a system of differential equations, and we numerically solve this system to obtain a solution of the optimization problem. This approach differs fundamentally from traditional central path methods, where solutions are obtained by solving a sequence of optimization problems at each step. Instead, we derive the entire solution path as the unique trajectory of a differential equation, providing a continuous and systematic way to reach the optimal solution without requiring iterative optimization at each stage.
@@
Related work
The homotopy method presented in this paper is deeply inspired by the principles of numerical algebraic geometry, as developed by Sommese and Wampler [28] (see [8, 29] for an overview). In this field, the primary goal is to solve a system of polynomial equations,
$p_{1}(\mathbf{z})=0,\ldots,p_{k}(\mathbf{z})=0,$
for multiple variables $\mathbf{z}\in\mathbb{C}^{n}$ with a finite number of solutions. The approach begins by constructing an initial system of polynomial equations,
$g_{1}(\mathbf{z})=0,\ldots,g_{k}(\mathbf{z})=0,$
whose solutions are trivial, such as the unit polynomials $g_{i}(\mathbf{z})=z_{i}^{d}$. The solutions of this system are then continuously tracked through the homotopy
$(1-t)g_{i}(\mathbf{z})+tp_{i}(\mathbf{z})=0,\quad\text{for every }i\in\{1,\ldots,k\}.$
In essence, the homotopy method involves gradually varying the parameter $t$ from 00 to $1$, while following the solution path by solving an associated differential equation.
In [9, 11], a homotopy approach was developed to solve semidefinite problems using the standard barrier-function relaxation. This method transforms problems lacking strict feasibility into a homotopy of problems that possess strict feasibility. However, these homotopies do not originate from a trivial initial optimization problem.
On the other hand, [14, 15] introduced the homotopy method for linear programs, aiming to transform a trivial problem into the target problem. In this approach, the homotopy is applied to the objective functions.
Furthermore, homotopic approach have also been applied in other contexts of linear and non-linear optimization [30, 12, 32].
Main contributions
To the best of our knowledge, our approach is the first to provide a general homotopy on feasible sets of convex optimization, connecting a trivial with a very general problem, and with proven convergence under mild conditions (see Theorem 6 for the main result of this paper). We prove that our approach applies successfully in the following cases:
@@
Overview
This paper is structured as follows: In Section 2 we introduce the notion of a homotopy between optimization problems, and then show under which conditions the induced differential equation of optimal solutions has a unique solution. In Section 3 we apply our homotopy method to two optimization classes: hyperbolic programs (which include semidefinite programs) and convex optimization problems with a single constraint function. In Section 4 we provide some numerical examples that compare our method to existing methods. In Section 5 we present open questions and further directions. Finally, Appendix A contains the proof of boundedness of rigidly convex sets defined by elementary symmetric polynomials, used in Section 4.
@@ 2.1 Homotopies of Convex Sets and Problem Statement
We do this by continuously deforming the feasible set from an easy one into $C$, and keeping track of an optimal solution of the corresponding optimization problems along the way. Note that this construction can be generalized to concave functions $f$, by introducing a homotopy of objective functions. In Remark 7 we elaborate on this more general case.
To be able to define and work with the homotopies, we will assume that the feasible sets along the homotopy (excluding the target feasible set) come with a description as follows:
@@
such that $p_{t}$ provides a smooth description of $C_{t}$ for all $t\in[0,1)$.
@@
Note that the descriptions of $C_{t}$ only needs to be smooth for $t<1$. Although the target feasible set $C_{1}$ does not need to be smooth (i.e. it has it has distinct vertices or kinks like in Fig. 1), there exist homotopies of smooth descriptions with $p_{1}$ providing a description of $C_{1}$.
@@
But under a certain continuity condition on the family $\left(C_{t}\right)_{t\in[0,1]}$, every homotopy of smooth descriptions is indeed regular. Assume, for example, that
$B:=\bigcup_{t\in[0,1)}\partial C_{t}\times\{t\}\subseteq\mathbb{R}^{n+1}$
is an $n$-dimensional manifold (with boundary $\partial C_{0}$). Then at each point $(\mathbf{x},t)\in B$ the gradient of $p$ is nonzero, already in space direction, since $p_{t}$ is a smooth description of $C_{t}$. Thus the implicit function theorem implies that the zero set of $p$ is the graph of a function in $t$ and $n-1$ of the space variables, locally around $(\mathbf{x},t)$. But since $B$ is an $n$-dimensional manifold contained in this zero set/graph, it coincides with it locally.
@@
This implies that $\mathbf{a}_{t}\in\partial C_{t}$ for every $t<1$. $\mathbf{a}_{1}\in\partial C_{1}$ follows by continuity of $t\mapsto\mathbf{a}_{t}$ and $t\mapsto\mathbf{p}_{t}(\mathbf{x})$.
@@ 2.2 The Differential Equation for Stationary Points
    $w_{i}=0$. Then $v_{i}w_{k}=0$. Since $w_{k}\neq 0$, we have that $v_{i}=0$ which implies that $v_{i}=\lambda w_{i}$.
@@ 2.3 Existence of a Unique Path of Optimal Solutions
Let $\left(C_{t}\right)_{t\in[0,1]}$ be a family of compact convex sets, all with non-empty interior in $\mathbb{R}^{n}$. Let $p$ be a homotopy of smooth descriptions satisfying the following properties:
@@
Remark 7.
Note that Theorem 6 remains valid even when the problem is generalized to allow:
  1. (i)
    a homotopy of objective functions.
  2. (ii)
    concave functions satisfying $\nabla f_{t}\neq 0$ on $C_{t}$, ensuring that the optimal point lies on the boundary.
Regarding (i), the construction of the differential equation follows the same process as in Section 2.2, with the only modifications being the replacement of $f$ by $f_{t}$ and the inclusion of its non-trivial time derivatives in $m(\mathbf{a}_{t},t)$. Since the invertibility of $K(\mathbf{x},t)$ remains unaffected by the choice of a homotopy for $f$, the result in Theorem 6 follows immediately.
Regarding (ii), the only remaining step is to show that $K(\mathbf{a}_{t},t)$ is invertible. This follows from a slight modification of the proof of Lemma 5. Since $\nabla f_{t}\neq 0$ on $C_{t}$, we have $\mathbf{a}_{t}\in\partial C_{t}$. By the concavity of $f_{t}$, it follows that
$\nabla f_{t}(\mathbf{a}_{t})^{t}(\mathbf{x}-\mathbf{a}_{t})\leqslant 0,$
We now prove that for every $\mathbf{v}\in\mathbb{R}^{n}$, we have
Assuming $\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})=0$ implies $\mathbf{v}^{t}\nabla f_{t}(\mathbf{a}_{t})=0$. Moreover, since $\lambda_{t}<0$ and $f_{t}$ is strictly concave, we obtain
$\mathbf{v}^{t}A(\mathbf{a}_{t},t)\mathbf{v}<0,$
where
For every $i\in\{1,\ldots,k-1,k+1,\ldots,n\}$, we rewrite
$\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}=\big{(}A(\mathbf{a}_{t},t)\mathbf{v}\big{)}_{k}\partial_{i}p_{t}(\mathbf{a}_{t})-\big{(}A(\mathbf{a}_{t},t)\mathbf{v}\big{)}_{i}\partial_{k}p_{t}(\mathbf{a}_{t}),$
where we used the relation $\partial_{i}f(\mathbf{a}_{t})=\lambda_{t}\partial_{i}p_{t}(\mathbf{a}_{t})$.
By Lemma 4, $\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}=0$ for all $i\in\{1,\ldots,n-1\}$ if and only if $A(\mathbf{a}_{t},t)\mathbf{v}$ is parallel to $\nabla p_{t}(\mathbf{a}_{t},t)$. However, since we have
$\mathbf{v}^{t}A(\mathbf{a}_{t},t)\mathbf{v}<0\quad\text{ and }\quad\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})=0,$
this leads to a contradiction.
@@ 3.1 Semidefinite and Hyperbolic Programming
Generalizing spectrahedra to rigidly convex sets (see Definition 8) leads to the notion of hyperbolic programming.
A hyperbolic program is a linear optimization problem over an affine slice of a hyperbolicity cone. Specifically, it is given by
$\max_{\mathbf{x}\in\mathbb{R}^{m}}$ $f(\mathbf{x})$
subject to $M\mathbf{x}=\mathbf{b},$
$\mathbf{x}\in\Lambda_{++}(h),$
where $\Lambda_{++}(h)$ denotes the hyperbolicity cone associated with a hyperbolic polynomial $h$ (see Section 4.1 for further details).
For a non-trivial affine slice, i.e., $M\neq 0$, hyperbolic programs can be reformulated as
$\max_{\mathbf{x}\in\mathbb{R}^{n}}$ $f(\mathbf{x})$
subject to $\mathbf{x}\in\mathcal{R}(p),$
where $\mathcal{R}(p)$ is the rigidly convex set in a smaller space, defined by a real-zero polynomial $p$ (see Definition 8).777 Assume there exists a vector $\mathbf{e}$ in the interior of $\Lambda_{++}(h)$ such that $M\mathbf{e}=\mathbf{b}$. If $\ker(M)$ has dimension $n=m-1$, we can rotate and rescale the space so that $\mathbf{e}=\mathbf{e}_{1}$. This transformation gives the real zero polynomial $p(\mathbf{x})=h(1,x_{1},\dots,x_{n}),$ with $\mathcal{R}(p)$ corresponding to $\Lambda_{++}(h)\cap\{x_{0}=1\}$ (see [20, Section 2.3] and [19, Definition 2.3.4]). If $\ker(M)$ has smaller dimension, this process can be iterated to obtain the formulation.
This class of sets generalizes semidefinite programming over spectrahedra (see Example 10).
Throughout this paper, we assume that the constraint $M\mathbf{x}=\mathbf{b}$ is non-trivial; otherwise, the feasible set would be $\Lambda_{++}(h)$, which is unbounded. This allows us to consider the hyperbolic program as an optimization problem over a rigidly convex set. Note that a similar situation occurs in semidefinite programming when $I_{s}$ is replaced by $\mathbf{0}$ in $A(\mathbf{x})$.
Rigidly convex sets arise from so-called real zero polynomials and share many properties with spectrahedra. The generalized Lax conjecture actually postulates that every rigidly convex set is a spectrahedron. While this statement is true for $n=2$, it is open in higher dimensions [13].
Note that our homotopy approach is useful beyond the established methods for solving semidefinite programs:
  • $\triangleright$
    It is not clear whether each rigidly convex set is actually a spectrahedron, so we solve a potentially larger class of problems. It seems that no efficient general methods to solve hyperbolic programs are available yet.
  • $\triangleright$
    Even if a set is known to be a spectrahedron, it might be hard to find an explicit linear matrix inequality defining it.
  • $\triangleright$
    Even if the linear matrix inequality is known, it may be much larger than the degree of the corresponding real-zero polynomial, while we work directly with the polynomial. This behavior is on the one hand generic [25], but there are also explicit examples with a proven super-polynomial overhead when described as spectrahedra [24].
@@ 3.1.1 Rigidly Convex Sets
In the following, we define and examine the notion of rigidly convex sets. Moreover, we review a procedure to approximate every rigidly convex set by a smooth version thereof. Most of the content in this subsection is also part of [19, Section 2].
@@
Let $A(\mathbf{x})=I_{s}+x_{1}A_{1}+\cdots+x_{n}A_{n}$ be a linear matrix function, and define the polynomial
@@
Note that multiplicities $k\geq 2$ arise when the spectrahedron can be written as the intersection of other spectrahedra. If $C(\mathbf{x})$ is the linear matrix inequality generated by matrices $A_{i}\oplus B_{i}$, then
$\mathcal{S}(C)=\mathcal{S}(A)\cap\mathcal{S}(B).$
Consequently, the corresponding real-zero polynomial satisfies
$p(\mathbf{x})=\det(C(\mathbf{x}))=\det(A(\mathbf{x}))\cdot\det(B(\mathbf{x})).$
If $\mathbf{x}\in\partial\mathcal{S}(A)$ and $\mathbf{x}\in\partial\mathcal{S}(B)$, then $s=1$ is a root of $s\mapsto p_{\mathbf{x}}(s)$ with multiplicity at least $k\geq 2$, reflecting the kink on the boundary of $\mathcal{S}(C)$ at $\mathbf{x}$.
@@ 3.1.2 Smoothing of Real Zero Polynomials
We introduce the smoothing operator $\mathcal{F}_{\varepsilon}$ on the space of real-zero polynomials, which transforms each rigidly convex set into a smooth rigidly convex set that approximates the original. This smoothing procedure was initially developed for hyperbolic polynomials [23, 27] (see also [19] for an overview). Since we are working in the affine setting, we introduce the smoothing operator directly at this level.
For this purpose, we introduce the Renegar derivative.
@@
Definition 12 (Smoothing of Real-Zero Polynomials).
Let $p\colon\mathbb{R}^{n}\to\mathbb{R}$ be a polynomial. We define the linear operator $\mathcal{F}_{\varepsilon}$ as
$\mathcal{F}_{\varepsilon}[p]\coloneqq p+\varepsilon\partial_{R}p,$
Applied to real-zero polynomials, we denote it by the $\varepsilon$-smoothed version of $p$.
@@
    If $\lambda$ is a root of multiplicity $k\geqslant 2$ of $p_{\mathbf{x}}$, then $\lambda$ is a root of multiplicity $k-1$ of $\mathcal{F}_{\varepsilon}[p]_{\mathbf{x}}$. All other roots of $\mathcal{F}_{\varepsilon}[p]_{\mathbf{x}}$ have multiplicity $1$.
@@
Remark 14.
By Lemma 13, the $m$-fold composition of the smoothing operator
$\mathcal{F}_{\varepsilon}^{m}(p)=p+\sum_{i=1}^{m}\varepsilon^{i}\partial_{R}^{i}p$
reduces the roots of $p_{\mathbf{x}}$ with multiplicity $k\geqslant m+1$ to roots with multiplicity $k-m$. All other roots of $\mathcal{F}_{\varepsilon}^{m}[p]_{\mathbf{x}}$ have multiplicity $1$.
Therefore, every rigidly-convex set $\mathcal{R}(p)$ can be smoothened by applying the smoothing operator $\mathcal{F}_{\varepsilon}$ sufficiently often. Moreover, if $\mathcal{R}(p)$ is bounded, the smoothed version also remains bounded.